# FAST Polarization Catalog of FRB 20240114A

### 21 Mar 2026 v1

@@ IV.1 Parameter Distributions
Linear/Circular Polarization Fraction (L/I, V/I, lower left panel): A significant fraction of bursts exhibit a high degree of linear polarization ( 76%), which is a common feature among repeating FRBs and suggests a magnetically ordered emission region. The distribution of the absolute circular polarization fraction is peaked at low values, with a majority of bursts below 40%. A tail of bursts extends to higher values, suggesting diverse emission properties or propagation effects within the sample.
@@ IV.2 Parameter Correlations
Figures 2-5 explore the relationships between several key parameters to further assess data integrity and identify global trends.
RM evolution (Figure 2 panel a): A plot of RM as a function of the topocentric arrival time (MJD) reveals the remarkable temporal evolution of this parameter. The data clearly show an initial period of relative stability followed by a significant linear decay, which will be the focus of a detailed analysis in a subsequent paper.
Polarization Fraction Distribution evolution (Figure 3): The top panel shows the observed burst number (with S/N 20) as a function of time, indicating significant variations in activity level. The middle panel quantifies the similarity between the two-dimensional (linear vs. circular) polarization fraction distributions of cumulative bursts in consecutive observing sessions with different time-scales using the Energy Distance metric. A larger Energy Distance corresponds to lower similarity between the two distributions. The notable peaks in this metric are found to correlate with abrupt, large changes in the number of detected bursts (as seen in the top panel), rather than signifying a fundamental change in the shape of the underlying polarization distribution itself. This conclusion is visually and statistically confirmed by the semicircle diagram in the bottom panel. Across the entire 16-month span, the dense locus of points consistently resides near the high-linear-polarization, low-circular-polarization end of the theoretical limit (the semicircle boundary). The core shape and concentration of this two-dimensional distribution show no systematic evolution correlated with the drastic RM change or the variable burst rate.
RM versus DM (Figure 4 left panel): No strong correlation is observed between RM and DM across the entire dataset (Pearson correlation coefficient r=-0.023). This lack of correlation suggests that the magneto-ionic material responsible for the Faraday rotation is not co-located with the dominant source of dispersion, implying a complex, multi-component environment.
@@
Polarization Fraction versus S/N (Figure 5): The linear and circular polarization fractions are plotted against the burst S/N (left and middle panel). As expected, the measurement uncertainty increases at low S/N. The absence of any spurious trends at high S/N confirms the robustness of our polarization calibration and debiasing procedures. The right panel shows the relationship between the linear (L/I) and circular (V/I) polarization fractions. A theoretical semi-circle with a radius of 100%, centered on the origin and extending in the positive L/I direction, is overlaid. This curve represents the physical limit for fully polarized emission. The points are color-coded by the burst S/N, and number-density contours are added to illustrate the distribution of the large dataset. Two key features are evident: (1) The vast majority of data points, especially those with high number density, cluster near the intersection of the semi-circle and the L/I axis (i.e., high linear and low circular polarization). (2) Points lying outside the theoretical semi-circle are predominantly those with low S/N. This is fully consistent with expectations, as measurement noise can spuriously push low-S/N points beyond the physical boundary, further validating our error estimation and the overall quality of the polarimetric measurements.
@@
In addition to the above parameters, we also calculated the proportion of bursts exhibiting significant circular polarization (degree of circular polarization, DOC 10%) in our sample. For FRB 20240114A, 1,157 out of 17,356 detected bursts met this criterion, yielding a proportion of 6.67% (see Figure 2 panel f). This measurement was examined within the context of the tentative anti-correlation between the significant circular polarization fraction and the absolute value of the source’s rotation measure () suggested for some repeaters. A power-law fit to our updated sample yields an index of . With this fit, the associated uncertainty envelope encompasses the data points for FRB 20190520B and FRB 20121102A (Zhang et al., 2023; Feng et al., 2022b), while those for FRB 20220912A and FRB 20201124A (Zhang et al., 2023) lie significantly outside of it. Should an underlying physical correlation exist, the fact that only a subset of sources aligns with the best-fit trend may point to intrinsic diversity within the repeating FRB population. Sources following the relation might originate in environments where the observed circular polarization is dominantly shaped by the same large-scale, ordered magneto-ionic medium responsible for the Faraday rotation (Niu et al., 2022; Anna-Thomas et al., 2023). In contrast, the significant deviation of other sources could imply that their polarization properties are governed by additional or distinct physical processes. For instance, complex local Faraday screens, burst emission geometries that favor strong intrinsic circular polarization, or magnetospheric propagation effects could decouple the observed circular polarization from the bulk RM in these objects (Hilmarsson et al., 2021; Xu et al., 2022; Mckinven et al., 2023). This dichotomy underscores that any universal relationship between circular polarization and RM is likely nuanced and modulated by source-specific conditions. It should be noted that this alignment with the trend for FRB 20240114A does not necessarily imply that the generation mechanism of circular polarization is identical to that of sources such as FRB 20121102A, given that no definitive signatures of Faraday conversion (systematic wavelength-dependent variations in circular polarization) were detected in our observations, which is consistent with the result of Uttarkar et al. (2026). Therefore, the circular polarization of FRB 20240114A may still be attributed to the intrinsic emission mechanism itself, rather than being dominated by propagation effects.
@@
In summary, the power-law fit reveals a shallower index for the DOC– relation, which is consistent with only a subset of repeating FRBs. This suggests that any such correlation is not universal but is likely modulated by source-specific local environments, as evidenced by the diverse RM behaviors predominantly attributed to proximate magneto-ionic screens. The dichotomy between sources that follow and those that deviate from the trend points to intrinsic diversity in either the emission mechanism or the magneto-ionic properties of the immediate vicinities of repeaters (Yang et al., 2023; Uttarkar et al., 2024). Further observations and larger samples are crucial to disentangle these physical effects.
@@
Having characterized the parameter distributions and correlations for FRB 20240114A, we next examine the frequency-dependent depolarization to explore the burst’s magneto-ionic environment. A random magnetic field component within the Faraday rotating screen would cause frequency-dependent depolarization, characterized by a decrease in the observed linear polarization fraction (L/I) towards lower frequencies. Depolarization due to the multipath propagation is given by (Feng et al., 2022a)
@@
where is the fractional reduction in the linear polarization amplitude, is the standard deviation of RM and is the wavelength. To test for this effect, we selected 18 bright bursts with high signal-to-noise ratio and broad bandwidth. The dynamic spectra and polarization profiles of these 18 bursts are shown in Figure 6. Each bursts’ band was subdivided into several frequency sub-bands. We independently performed RM synthesis and measured L/I within each sub-band. As shown in Figure 6, the derived L/I values for all sub-bands across all selected bursts cluster near 100%, with no systematic trend as a function of frequency. To better illustrate whether the L/I varies significantly with frequency, we plotted the L/I distribution of slices for each burst (left panel of Figure 6). To avoid overlap, the bursts were randomly divided into three groups. If the L/I distribution of a burst exhibits significant broadening, it would indicate a pronounced frequency evolution of the L/I, whereas a concentrated distribution implies that the L/I is nearly independent of frequency. The results show that the L/I distributions of all 18 bursts are relatively concentrated, indicating that the L/I remains nearly constant across frequency. This absence of significant depolarization allows us to place an upper limit on any random magnetic field component (often quantified as ) within the main Faraday screen. Consequently, the RM values reported in our catalog for these and similar bursts can be reliably interpreted as representing a largely ordered magnetic field along the line of sight, with minimal internal Faraday dispersion.
The overall properties and distributions presented here confirm the high quality and internal consistency of the catalog. The complex structures and correlations visible in these global views underscore the scientific value of this large, uniform dataset for probing the dynamic nature of FRB 20240114A.
@@ V.1 The Origin of RM Evolution
Our session-based analysis (Figure 7) starkly contrasts the temporal behaviors of DM and RM. The DM remains stable across all six activity-defined sessions, with its distribution peak and width showing negligible changes (Figure 7, panel e). This indicates a constant total electron column density along the line-of-sight over 16 months. In dramatic contrast, the RM undergoes a highly significant, monotonic decay of approximately 200 . The shift in the RM distribution peak between sessions is visually clear (Figure 7, panel e). The Monte Carlo simulations confirm that the observed drop, particularly during session V, is significant at a level exceeding 121 compared to the null hypothesis of a constant magnetic field (red dashed line in Figure 7, panel c).
The evolution of the line-of-sight magnetic field component can be quantitatively estimated from the changes in RM and DM, which is given by , where and are the differences in RM and DM between different sessions. Applying this to the significant RM decay observed between activity sessions, we find that itself exhibits a corresponding decrease (Figure 7, panel d). The uncertainties of are derived from error propagation of the measurement errors in RM and DM. This direct estimate demonstrates that the evolution of is consistent with the RM decay, and is not in contradiction with it. The stable standard deviation of DM over time (Figure 7, panel a) argues against the injection of new, clumpy plasma, supporting the interpretation that the changing RM and stem from the evolution of an existing magneto-ionic environment.
@@ V.2 Complex and Variable Emission Geometry
The polarization position angle evolution of FRB 20240114A also exhibits this dynamic characteristic. Unlike FRB 20180916B, which shows remarkably stable burst-to-burst PAs indicative of a highly stable magnetic field geometry at the emission site (Bethapudi et al., 2025), FRB 20240114A exhibits large burst-to-burst PA variation (Figure 8). The broad distribution of PA0, even for bursts with stable intra-burst PA profiles (blue histogram in Figure 8), and the significant scatter of the session-averaged PA0 (red error bars in Figure 8, left) suggest that the bursts originate from a spatially complex or temporally fast-evolving magnetospheric environment. The lack of a stable, globally preferred PA challenges models where emission is strictly confined to a fixed magnetic pole. The combination of a systematically evolving RM (probing the large-scale, intervening screen) and a highly variable PA (probing the immediate emission geometry) suggests that both the large-scale environment and the small-scale emission region are dynamic, albeit likely on different timescales.
Interestingly, independent observations of FRB 20240114A with the Parkes ultra-wideband low (UWL) receiver have revealed a multi-peak distribution in the polarization position angle, with excesses near -17.39 degree and 64.8 degree (Uttarkar et al., 2026). This contrasts with the PA distribution observed in our FAST data, suggesting that the apparent PA distribution may depend on the observed frequency band or the specific activity epoch. Such frequency-dependent PA behavior would be expected in a dynamically evolving magnetosphere, where different emission heights or regions could be probed at different frequencies. However, the effect appears to be relatively weak, as both datasets consistently point to a highly variable emission geometry without a single preferred direction.

### 18 Apr 2026 v2

@@ IV.1 Parameter Distributions
Linear/Circular Polarization Fraction (, , lower left panel): The linear polarization fractions are systematically high, with a 3 lower bound of 76%, a behavior commonly seen among repeating FRBs and suggestive of a magnetically ordered emission region. The distribution of the absolute circular polarization fraction is peaked at low values, with a majority of bursts below 40%. A tail of bursts extends to higher values, suggesting diverse emission properties or propagation effects within the sample.
@@ IV.2 Parameter Correlations
Figures 2-5 explore the relationships among several key parameters to assess the overall consistency of the measurements and to identify global trends in the burst sample.
RM evolution (Figure 2): Among the parameters shown in Figure 2, RM is the only one that exhibits a pronounced long-term evolution, consistent with the trend previously noted by Uttarkar et al. (2026b). As shown in panel (a), RM as a function of topocentric arrival time displays an initial period of relative stability followed by a significant linear decay. In contrast, the corresponding time evolutions of DM, total polarization fraction (), linear polarization fraction (), and circular polarization fraction () shown in panels (b)–(e) do not display a similarly distinct transition, and instead remain comparatively stable or scattered without a clear monotonic trend. This RM evolution will be the focus of a more detailed analysis in a subsequent paper.
Polarization Fraction Distribution evolution (Figure 3): The top panel shows the observed burst number (with S/N 20) as a function of time, indicating significant variations in activity level. The middle panel quantifies the similarity between the two-dimensional (linear vs. circular) polarization fraction distributions of cumulative bursts in consecutive observing sessions with different time-scales using the Energy Distance metric (see Appendix D). A larger Energy Distance corresponds to lower similarity between the two distributions. The notable peaks in this metric are found to correlate with abrupt, large changes in the number of detected bursts (as seen in the top panel), rather than signifying a fundamental change in the shape of the underlying polarization distribution itself. This conclusion is visually and statistically confirmed by the semicircle diagram in the bottom panel. Across the entire 16-month span, the dense locus of points consistently resides near the high-linear-polarization, low-circular-polarization end of the theoretical limit (the semicircle boundary). The core shape and concentration of this two-dimensional distribution show no systematic evolution correlated with the drastic RM change or the variable burst rate.
RM versus DM (Figure 4 left panel): No strong correlation is observed between RM and DM across the entire dataset (Pearson correlation coefficient r=-0.023). This lack of correlation suggests that the RM variability is not controlled by changes in the total electron column density alone. Instead, it is consistent with a scenario in which the Faraday rotation is dominated by variations in the line-of-sight magnetic field component, including possible changes in its strength or orientation, within a complex magneto-ionic environment.
@@
Polarization Fraction versus S/N (Figure 5): The linear and circular polarization fractions are plotted against the burst S/N (left and middle panel). As expected, the measurement uncertainty increases at low S/N. The absence of any spurious trends at high S/N confirms the robustness of our polarization calibration and debiasing procedures (see Appendix B). The right panel shows the relationship between the linear () and circular () polarization fractions. A theoretical semi-circle with a radius of 100%, centered on the origin and extending in the positive direction, is overlaid. This curve represents the physical limit for fully polarized emission. The points are color-coded by the burst S/N, and number-density contours are added to illustrate the distribution of the large dataset. Two key features are evident: (1) The vast majority of data points, especially those with high number density, cluster near the intersection of the semi-circle and the axis (i.e., high linear and low circular polarization). (2) Points lying outside the theoretical semi-circle are predominantly those with low S/N. This is fully consistent with expectations, as measurement noise can spuriously push low-S/N points beyond the physical boundary, further validating our error estimation and the overall quality of the polarimetric measurements.
@@
In addition to the above parameters, we also calculated the proportion of bursts exhibiting significant circular polarization (degree of circular polarization, 10%) in our sample. For FRB 20240114A, 1,157 out of 17,356 detected bursts met this criterion, yielding a proportion of 6.67% (see Figure 2 panel f). This measurement was examined within the context of the tentative anti-correlation between the significant circular polarization fraction and the absolute value of the source’s rotation measure () suggested by Zhang et al. (2023). A power-law fit to our updated sample yields an index of . With this fit, the associated uncertainty envelope encompasses the data points for FRB 20190520B and FRB 20121102A (Zhang et al., 2023; Feng et al., 2022b), while those for FRB 20220912A and FRB 20201124A (Zhang et al., 2023) lie significantly outside of it. Should an underlying physical correlation exist, the fact that only a subset of sources aligns with the best-fit trend may point to intrinsic diversity within the repeating FRB population. Sources following the relation might originate in environments where the observed circular polarization is dominantly shaped by the same large-scale, ordered magneto-ionic medium responsible for the Faraday rotation (Niu et al., 2022; Anna-Thomas et al., 2023a). In contrast, the significant deviation of other sources could imply that their polarization properties are governed by additional or distinct physical processes. For instance, complex local Faraday screens, burst emission geometries that favor strong intrinsic circular polarization, or magnetospheric propagation effects could decouple the observed circular polarization from the bulk RM in these objects (Hilmarsson et al., 2021; Xu et al., 2022; Mckinven et al., 2023b). This dichotomy underscores that any universal relationship between circular polarization and RM is likely nuanced and modulated by source-specific conditions. It should be noted that this alignment with the trend for FRB 20240114A does not necessarily imply that the generation mechanism of circular polarization is identical to that of sources such as FRB 20121102A. Therefore, the circular polarization of FRB 20240114A may still be attributed primarily to the intrinsic emission mechanism itself, rather than being dominated by propagation effects.
@@
In summary, the power-law fit reveals a shallower index for the – relation, which is consistent with only a subset of repeating FRBs. This suggests that any such correlation is not universal but is likely modulated by source-specific local environments, as evidenced by the diverse RM behaviors predominantly attributed to proximate magneto-ionic screens. The dichotomy between sources that follow and those that deviate from the trend points to intrinsic diversity in either the emission mechanism or the magneto-ionic properties of the immediate vicinities of repeaters (Yang et al., 2023; Uttarkar et al., 2024). Continued observations of FRB 20240114A, together with comparisons across a larger sample of repeating FRB sources, will be crucial for disentangling these physical effects.
@@
Having characterized the parameter distributions and correlations for FRB 20240114A, we next examine the frequency-dependent depolarization to explore the burst’s magneto-ionic environment. A random magnetic field component within the Faraday rotating screen would cause frequency-dependent depolarization, characterized by a decrease in the observed linear polarization fraction () towards lower frequencies. In the classical Burn (1966) framework for depolarization by Faraday dispersion, the fractional reduction in linear polarization can be written as
@@
where is the fractional reduction in the linear polarization amplitude, is the standard deviation of RM and is the wavelength. This formalism has since been applied to FRBs by Feng et al. (2022a), Uttarkar et al. (2026a), and Pandhi et al. (2026). To test for this effect, we selected 18 bright bursts with high signal-to-noise ratio and broad bandwidth. The dynamic spectra and polarization profiles of these 18 bursts are shown in Figure 6. Each bursts’ band was subdivided into several frequency sub-bands. We independently performed RM synthesis and measured within each sub-band. As shown in Figure 6, the derived values for all sub-bands across all selected bursts cluster near 100%, with no systematic trend as a function of frequency. To better illustrate whether the varies significantly with frequency, we plotted the distribution of slices for each burst (left panel of Figure 6). To avoid overlap, the bursts were randomly divided into three groups. If the distribution of a burst exhibits significant broadening, it would indicate a pronounced frequency evolution of the , whereas a concentrated distribution implies that the is nearly independent of frequency. The results show that the distributions of all 18 bursts are relatively concentrated, indicating that the remains nearly constant across frequency. This absence of significant depolarization allows us to place an upper limit on any random magnetic field component (often quantified as ) within the main Faraday screen. Based on our Bayesian analysis of 18 bursts, we derive a 95% upper limit of 0.51 rad m-2.This suggests that the main Faraday screen is largely ordered and exhibits minimal internal Faraday dispersion, at least for the selected bursts analysed here. A broader population-level test based on burst central frequency, bandwidth, and would be valuable, but is not yet sufficiently robust for the present data set because our current CDF-based bandwidth measurements do not always yield reliable central-frequency estimates, particularly for bursts truncated by the band edges. In addition, for multi-component bursts, differences in among distinct temporal or spectral components may contaminate a depolarization search by introducing apparent frequency-dependent behavior that does not arise from a single, uniform Faraday screen.
We additionally searched for signatures of Faraday conversion in bursts with sufficient S/N and frequency coverage. Specifically, we examined whether the circular polarization fraction exhibits any systematic frequency- or wavelength-dependent evolution within individual bursts. We do not find any clear or definitive evidence for such behavior in the present sample. This is also consistent with the result of Uttarkar et al. (2026b).
@@ V.1 The Origin of RM Evolution
Our session-based analysis (Figure 7) starkly contrasts the temporal behaviors of DM and RM. The DM remains stable across all six activity-defined sessions, with its distribution peak and width showing negligible changes (Figure 7, panel e). This indicates that no strong secular DM evolution is detected across the six activity-defined sessions. To quantify the additional burst-to-burst DM scatter beyond the formal fitting uncertainties, we compared the burst-level DMs with the corresponding daily-average reference DM values and wrote the observed daily scatter as , where is the formal measurement contribution, represents the morphology-related excess variance, and denotes any intrinsic day-scale DM variation. For our sample, we obtain and , implying . Under the practical assumption that , this yields an upper limit of . We therefore interpret the absence of an obvious long-term DM trend as a non-detection above the burst-level variance floor, rather than as evidence for perfectly constant DM. This interpretation is also supported by the recent FAST study of Feng et al. (2026), who argued that part of the apparent burst-to-burst DM scatter in repeating FRBs can be morphology-induced (“pseudo DM”) rather than reflecting true propagation changes. Our result for FRB 20240114A is qualitatively consistent with this picture. In dramatic contrast, the RM undergoes a highly significant, monotonic decay of approximately 200 . The shift in the RM distribution peak between sessions is visually clear (Figure 7, panel e).
To quantify the significance of this decline, we performed a Monte Carlo test using the burst-level RM measurements from sessions I–IV as the null hypothesis of a stable RM distribution. Specifically, we combined all RM values from these four sessions and used their sample mean and standard deviation to define the parent distribution. We then generated 10,000 mock samples, each containing the same number of bursts as observed in sessions V–VI, by drawing random values from this null distribution. For each mock sample, we calculated its mean RM and compared the resulting distribution of simulated means with the observed mean RM of sessions V–VI. The observed value lies far below the simulated distribution, and none of the 10,000 realizations reproduces such a low mean, implying p 10-4 under the null hypothesis of no RM evolution.
The evolution of the line-of-sight magnetic field component can be quantitatively estimated from the changes in RM and DM, which is given by , where and are the differences in RM and DM between different sessions. Applying this to the significant RM decay observed between activity sessions, we find that the nominal values of vary in a manner qualitatively consistent with the RM evolution (Figure 7, panel d). However, the uncertainties are large, and the current data do not provide statistically significant evidence for a departure from a constant temporal evolution of . The uncertainties of are derived from error propagation of the measurement errors in RM and DM. The stable standard deviation of DM over time (Figure 7, panel a) argues against the injection of new, clumpy plasma, supporting the interpretation that the changing RM and stem from the evolution of an existing magneto-ionic environment.
This inferred local line-of-sight magnetic field can also be compared with similar estimates discussed for other repeating FRBs. For FRB 20240114A, our session-based estimates span from values consistent with zero at early times to tens and up to at later epochs, with nominal values ranging from about -168 to +54 G (i.e., 0.05-0.17 mG in magnitude). This places FRB 20240114A below the more strongly magnetized mG-class environments inferred for FRB 20121102A (Michilli et al., 2018), FRB 20190520B (Anna-Thomas et al., 2023b), FRB 20190417A (Moroianu et al., 2026), and FRB 20220529A (Pandhi et al., 2026), and also somewhat below FRB 20240619D (Ould-Boukattine et al., 2026). In this sense, FRB 20240114A appears less extreme than highly magnetized systems, while still indicating a dynamically evolving magneto-ionic environment. More broadly, it is comparable to the class of repeaters for which RM and DM evolution imply a local environment that is clearly magnetized but not as extreme as the most strongly Faraday-active sources.
@@
A useful phenomenological comparison is FRB 20180916B, whose RM evolution also exhibits an initially relatively stable phase followed by a more secular change (Mckinven et al., 2023a). In that sense, FRB 20240114A does not appear to be an isolated case, but may belong to a subset of repeaters whose magneto-ionic environments evolve smoothly over month-to-year timescales rather than through purely stochastic fluctuations. At the same time, the detailed amplitudes, durations, and possible reversals of the RM evolution differ between sources, indicating diversity in the underlying environments and evolutionary drivers.
@@ V.2 Complex and Variable Emission Geometry
The polarization position angle evolution of FRB 20240114A also exhibits this dynamic characteristic. For each burst, we first determine its RM and derotate the Stokes and data accordingly. We then construct a time-resolved PA profile from the frequency-summed derotated Stokes parameters. The uncertainty of each PA sample, denoted as PAE, is estimated by propagating the off-pulse noise in and ; only on-pulse samples with PAE are retained. We define PA0 as the time average of the retained derotated PA samples for each burst. Unlike FRB 20180916B, which shows remarkably stable burst-to-burst PAs indicative of a highly stable magnetic field geometry at the emission site (Bethapudi et al., 2025), FRB 20240114A exhibits large burst-to-burst PA variation (Figure 8). Using the reduced- of a constant-PA fit to the derotated time-resolved PA profiles, we find that 67.2% of bursts satisfy . However, does not necessarily imply a prominent PA swing; if we additionally require the difference between the maximum and minimum PA within a burst to exceed 40∘, the fraction of bursts with clear intraburst PA swings is 17.6%. For bursts with variable PA profiles, PA0 should therefore be regarded as a representative value rather than a complete description of the full intraburst PA evolution. The broad distribution of PA0, even for bursts with stable intra-burst PA profiles (blue histogram in Figure 8), and the significant scatter of the session-averaged PA0 (red error bars in Figure 8, left) suggest that the bursts originate from a spatially complex or temporally fast-evolving magnetospheric environment. The lack of a stable, globally preferred PA challenges models where emission is strictly confined to a fixed magnetic pole. The combination of a systematically evolving RM (probing the large-scale, intervening screen) and a highly variable PA (probing the immediate emission geometry) suggests that both the large-scale environment and the small-scale emission region are dynamic, albeit likely on different timescales.
Interestingly, independent observations of FRB 20240114A with the Parkes ultra-wideband low (UWL) receiver have revealed a multi-peak distribution in the polarization position angle, with excesses near -17.39 degree and 64.8 degree (Uttarkar et al., 2026b). This contrasts with the PA distribution observed in our FAST data, suggesting that the apparent PA distribution may depend on the observed frequency band or the specific activity epoch. Such frequency-dependent PA behavior would be expected in a dynamically evolving magnetosphere, where different emission heights or regions could be probed at different frequencies. However, the effect appears to be relatively weak, as both datasets consistently point to a highly variable emission geometry without a single preferred direction.